OpenStudy (bananas1234):

If f(x) varies directly with x, and f(x) = 56 when x = 8, find the value of f(x) when x = 2 4 7 8 14

OpenStudy (bananas1234):

@Nnesha would it be possible for you to help me with one more?

Directrix (directrix):

y = k*x 56 = k * 8 What is the value of k?

OpenStudy (bananas1234):

k = 56/8, But what happens to the 2 that is listed in the equation?

Directrix (directrix):

You can read more about direct variation here, if interested: http://www.regentsprep.org/regents/math/algebra/AO4/Ldirect.htm

Directrix (directrix):

We have to get the equation first. After that, we deal with the 2.

OpenStudy (bananas1234):

ok

OpenStudy (bananas1234):

so 56/8 needs to be simplyfied

Directrix (directrix):

What is 56 divided by 8. We need that for the variation constant.

OpenStudy (bananas1234):

to 7?

OpenStudy (bananas1234):

so K=7

Directrix (directrix):

So, the equation is this: y = 7*x

Directrix (directrix):

If x = 2, then use the equation to get y. Post what you get.

OpenStudy (bananas1234):

y = 14

OpenStudy (bananas1234):

Then what do i do next?

Directrix (directrix):

That is it. That is the value of f(x) when x = 2 --> 14

OpenStudy (bananas1234):

ohh i see. Thank you!

Directrix (directrix):

You are welcome.

OpenStudy (bananas1234):

I need help with one more, should i open a new question?

Directrix (directrix):

Yes.

OpenStudy (bananas1234):

ok, thanks

Directrix (directrix):

You are welcome.